# State-Space Modeling in Long Sequence Processing: A Survey on Recurrence in the Transformer Era

### v1

@@ 4 Deep State-Space Models
Recent works on models that are intrinsically based on recurrent computations particularly emphasize the notion of (deep) State-space Model [90]. In particular, there exists a growing interest in exploiting the computational advantages of using multiple stacked instances of linear recurrences, whose dynamics is appropriately conditioned to avoid trivial explosive/vanishing dynamics. The development of such line of works can be traced back to the seminal work of [293] and [86] in which the authors propose methods to perform online function approximation. Then, the scientific literature of the last years transitioned from focusing on online function approximation to specifically designed (deep) State-Space Models [90, 88] and more advanced architectures exploiting them [85, 57], as we anticipate in Figure 15.
Online Function Approximation.   The basics of online function approximation, for what concerns the first works on this novel wave of state-space models, can be formalized as follows. Given a function of one variable defined on the half line $u\colon[0,+\infty)\to\mathbb{R}$, the problem of online approximation of such function is twofold: ($i$) for time instant $t\in[0,+\infty)$ find an approximation of $u$ until $t$, i.e., $u^{t}:=u|_{I_{t}}$ with $I_{t}:=(0,t)$ and ($ii$) have a method to update online such approximation. In order to formalize the concept of approximation, we need to have some notion of closeness, and hence we assume that the function that we want to approximate lives in some normed space. Moreover, the measure with respect to which we define the notion of integrability has a rather important role for computing an online approximation. In the following descriptions, we will mostly refer to the seminal works in [86], where the HiPPO model/theory is introduced, and [293], based on Legendre Memory Units (LMUs). In [86] the authors find that working with a normalized Lebesgue measure on $I_{t}$ (which is the standard choice in $\mathbb{R}^{n}$) has several advantages. A different choice is explored in LMU [293] that, in light of the theoretical formulation of the problem presented in [86], correspond to choosing a measure of density that is constant on a window $[t-\theta,t]$ of size $\theta$ just before the end point $t$ of the considered temporal instant. The other basic ingredient to consider in function approximation is the class of basis functions with which we want to perform such an approximation. In [86] the authors consider the case of translated and rescaled Legendre polynomials, $v^{t}_{n}$ for $n=1,2,\dots$, defined in $[0,t]$ by,
@@
[…]
@@
[…]
@@
$A_{ij}(t)=-\frac{1}{t}\begin{cases}\sqrt{(1+2i)(1+2j)}&\hbox{for$i>j$}\\ 1+i&\hbox{for $i=j$}\\ 0&\hbox{for $i<j$}\\ \end{cases},$ (39)
@@
where $\Phi$ is the transition function that depends on the precise recurrent architecture, $u_{t}$ is the input of the net at time $t$ and $c_{t-1}$ are the coefficients of the online approximation of the function $f_{t}=\pi(x_{t})$, where $\pi\colon\mathbb{R}^{d_{\text{s}}}\to\mathbb{R}$ is a projection of the state onto the real line (which is necessary, since these online approximation methods work on scalar functions). The leap from this hybrid model, where the state of the recurrence is enriched with an online approximation of the state itself, to Deep State-Space Models has been proposed in [90], in which Linear State Space Layers (LSSLs) are analyzed in comparison to other deep learning models that are used to process sequences, and in [88], where such models are refined to overcome computational limitations of the former model. In linear state space layers (Figure 13) the main idea is to use a linear continuous time expression to model the update of the state itself,
@@
The continuous time model described by Eq. (41) is typically discretized in order to be numerically implemented. Different discretization techniques can be applied, the more direct being the explicit (or forward) Euler method,151515 Technically this is a mixed scheme since the input is computed at the step $t+1$, however we call it explicit since it is so with respect to the state variable.
@@
This is what is commonly referred to as the convolutional form of the linear state space model in Eq. (44) – see Figure 14-top. Now, going back to LSSL [90], the matrix $A$ is represented with a suitable matrix factorization, i.e., $A=P(D+T^{-1}Q)$, with $D$, $P$ and $Q$ diagonal matrices and $T$ tridiagonal. The HiPPO matrix $A(t)$ defined in Eq. (39) admits such factorization (see Appendix E.2 of [90]). In this way, matrix $A$ is guaranteed to be quasiseparable, a property that is presented as desirable both for handling long dependencies and for having efficient matrix-vector multiplication. As a common practice in deep learning, several LSSLs can be stacked together, each layer receiving as input the output of the previous one. This is possible since the input and the output are of the same dimension. The main problem with this architecture is the computational cost; indeed (see [88]), it requires $\mathcal{O}({d_{\text{s}}}^{2}L)$ operations and it scales as $\mathcal{O}({d_{\text{s}}}L)$ for what concerns memory in order to compute the input-output mapping in Eq. (44). In order to overcome this precise limitation the S4 model [88] has been introduced to condition the structure of the matrix $A$. There is a large set of works that were published in the last few years along this line of research, and that, starting from S4, we describe in the following. Refer to Figure 15 for an overview.
@@
[…]
@@
From SISO to MIMO: S5.   The Simplified Structured State Space Sequence Model (S5) [271] is the first Deep SSM to be parameterized leveraging the concept of multiple input multiple output (MIMO) systems, for simplifying the architectural components and enhancing computations. This means that the full input vector $u\in\mathbb{R}^{d_{\text{in}}}$ is fed into a single bigger MIMO system of Eq. (48), instead of ${d_{\text{in}}}$ SISO scalar smaller subsystems, by stacking the matrices $A^{\tau},\,B^{\tau},\,C$, used in S4 and S4D. S5 inherits the S4D parametrization of the matrix $A$ (i.e., it is a diagonal matrix), while it can be discretized applying both bilinear (as done in S4, see Eq. (47)) and exact (as done in S4D, see Eq. (50)) discretizations. The MIMO structure and the diagonal parameterization allows for parallel computation of the individual output components via a parallel scan algorithm (see Appendix H of [271]). The parallel scan algorithm (see Figure 14-bottom) offers a way to parallelize the computations of a sequence of elements of a semigroup $(S,\bullet)$ generated by a recurrence relation of the the form $s_{i+1}=s_{i}\bullet c_{i}$, where $(c_{i})_{i=1}^{L}$ is a given sequence of elements of $S$161616 We recall that a semigroup consist of a set $S$ together with an associative operation $\bullet$. This approach can be directly applied to the computation of a linear recurrence of the form in Eq. (44) with the following choices,
@@
Figure 15: Timeline on the development of deep state-space models—Overview of the organization of Section 4.
SSMs in Language Modeling: H3.   The Hungry Hungry Hippo (H3) [55] model is a novel approach to leverage SSMs for language modeling, aiming to address the limitations of previous SSMs in tasks like Associative Recall and Induction Heads (see Section 7) compared to attention-based models. H3 draws inspiration from linear attention, which assumes a specific form for the similarity metric used in attention calculations. It stacks two discrete SSMs: one with a shift matrix (i.e., a local convolution) to remember past tokens and one with a diagonal matrix to retain state over the entire sequence. The key innovation lies in introducing gates (i.e., multiplicative interactions) between the outputs of these SSMs and projections of the input that, combined with the shift matrix, enable H3 to compare tokens across the sequence. The shift SSM identifies specific events (like the presence of a key token), while the diagonal SSM stores and retrieves associated information (like the corresponding value token). H3 has a time complexity of $\mathcal{O}(Llog(L))$ for a sequence of length $L$, making it asymptotically more efficient than traditional attention, which has a complexity of $\mathcal{O}(L^{2})$.
@@
where $C^{\tau}_{t}=C_{t}$, $D^{\tau}_{t}=D_{t}$ and $W_{\tau}\in\mathbb{R}^{1\times{d_{\text{in}}}}$. The model is structured in a MIMO manner (as S5, LRU and S6) and the dynamic matrix $A$ is initialized with $\lambda_{j}=-j$, for every $j=1,\dots,{d_{\text{s}}}$, ensuring that the eigenvalues lie in the negative halfplane. Since $\tau_{t}$ is time-varying, the eigenvalues of $A^{\tau}_{t}$ have an initialization that is input-dependent. The time-varying representation presents computational challenges, despite being more expressive. The authors provide an efficient implementation of the time-varying dynamics in Eq. (48), presenting a variation of the parallel scan and exploiting it both at inference and training time. S6 introduces an innovative way of pre-processing the input, called Mamba, which relies on both linear and non-linear maps. The input enters the recurrence through a linear projection, followed by a causal convolution. On the other hand, it passes through a linear projection followed by a SiLU nonlinearity before entering the gating function for post-processing. The gating function is inspired to previous models, i.e., H3 and GAU. An architecture close to Mamba is the Gated State Space (GSS) layer [189], again inspired by GAU. GSS resembles Mamba block but includes additional projections. The key difference is that GSS’s projection reduces the model dimension to decrease the state size of the SSM, whereas Mamba expands the model dimension to increase the state size.
@@
for $i=0,\dots,{d_{\text{in}}}-1$, where $h^{j}_{t}$ are implicit long convolution filters learned by shallow feed-forward neural networks and $R\in\mathbb{R}^{{d_{\text{in}}}\times{d_{\text{in}}}}$ is an output projection that mixes channels across the sequence length. This approach decouples filter length from parameter cost, providing advantages over explicit parameterization. The number of recurrent steps determines the complexity of the operator and they can be represented as a decomposition of data-controlled matrices. These matrices dynamically adapt based on the input data, similar to how attention mechanisms compute a weighted sum over input elements. Instead of explicitly computing the full data-controlled matrix, Hyena leverages fast convolution algorithms, particularly Fast Fourier Transform (FFT)-based convolutions, to achieve subquadratic time complexity. Moreover, unlike some models that restrict the receptive field, it allows for interactions between any elements in the sequence through its long convolutions, enabling it to capture long-range dependencies effectively. In subsequent work [186], some improvements to further enhance the efficiency of Long Convolution Sequence Models (LCSMs), including Hyena, have been introduced: the LaughingHyena distilling. It focuses specifically on improving the inference stage of these models, particularly in auto-regressive generation tasks. LaughingHyena distillation process aims to represent each convolutional filter of a pre-trained LCSM into an SSM, with the smallest state dimension such that it approximates the original filter without significant loss of accuracy. To achieve this goal, it utilizes a method called modal interpolation, which provides coefficients for the numerator and denominator of a rational function that minimizes the difference between the original filter and the approximating SSM transfer functions. Through these coefficients, it is possible to define the matrices $A$, $B$ and $C$ which characterize the SSM. This distillation procedure is then followed by two steps: ($i$) pre-filling and ($ii$) a recurrent update rule. Pre-filling involves computing the state to start generating new tokens when a length-$L$ prompt is fed to the model during auto-regressive generation, exploiting the denominator of the approximate transfer function. The recurrent update rule for the complex state is instead given by,
@@
Theoretical Foundations of SSMs.   Performances achieved by SSMs are remarkable, thus inspiring several researches to understand both their expressive capabilities and the connections to existing popular technologies (such as attention), with which they share many features but have been commonly developed in isolation. Orvieto et al. [212] theoretically show that combining MLPs with either real or complex linear diagonal recurrences (such as in S4, Mamba, etc.) enables highly precise approximation of regular causal sequence-to-sequence maps. The proof is based on the fact that the linear RNN provides a lossless encoding of the input sequence, and the MLP conducts non-linear processing on this encoding. While real diagonal linear recurrences are sufficient for achieving universality, employing complex eigenvalues near the unit disk, a strategy that has shown empirical success in S4, significantly improves the ability of recurrent model to store information. Cirone et al. [44] leverages tools from Rough Path Theory, and provide theoretical grounding for the fact the when random linear recurrences are enhanced with simple input-controlled transitions (selectivity mechanism), the hidden state is demonstrably a low-dimensional projection of a mathematical construct known as the signature of the input. This signature captures non-linear interactions between tokens across different timescales. Other recent works focus on the connections and differences between SSMs and other sequence processing models [101, 266], as long as to their links with control theory [2].

### v2

@@ 4 Deep State-Space Models
Recent works on models that are intrinsically based on recurrent computations particularly emphasize the notion of (deep) State-space Models [40]. In particular, there exists a growing interest in exploiting the computational advantages of using multiple stacked instances of linear recurrences, whose dynamics are appropriately conditioned to avoid trivial explosive/vanishing dynamics. The development of such lines of works can be traced back to the seminal work of [30] and [39], in which the authors propose methods to perform online function approximation. Then, the scientific literature of the last years transitioned from focusing on online function approximation to specifically designed (deep) State-Space Models [40, 25] and more advanced architectures exploiting them [146, 44], as we anticipate in Figure 15.
Online Function Approximation.   The basics of online function approximation, with regard to the first works on this novel wave of state-space models, can be formalized as follows. Given a function of one variable defined on the half line $u\colon[0,+\infty)\to\mathbb{R}$, the problem of online approximation of such function is twofold: ($i$) for time instant $t\in[0,+\infty)$ find an approximation of $u$ until $t$, i.e., $u^{t}:=u|_{I_{t}}$ with $I_{t}:=(0,t)$ and ($ii$) have a method to update online such approximation. In order to formalize the concept of approximation, we need to have some notion of closeness, and hence we assume that the function we want to approximate lies in some normed space. Moreover, the measure with respect to which we define the notion of integrability plays a rather important role in computing an online approximation. In the following descriptions, we will mostly refer to the seminal works in [39], where the HiPPO model/theory is introduced, and [30], based on Legendre Memory Units (LMUs). In [39] the authors find that working with a normalized Lebesgue measure on $I_{t}$ (which is the standard choice in $\mathbb{R}^{n}$) has several advantages. A different choice is explored in LMU [30] that, in light of the theoretical formulation of the problem presented in [39], correspond to choosing a measure of density that is constant on a window $t-\theta,t]$ of size $\theta$ just before the end point $t$ of the considered temporal instant. The other basic ingredient to consider in function approximation is the class of basis functions with which we want to perform such an approximation. In [39] the authors consider the case of translated and rescaled Legendre polynomials, $v^{t}_{n}$ for $n=1,2,\dots$, defined in $[0,t]$ by,
@@
where $(u^{t},v_{n}^{t})_{t}:=\int_{I_{t}}u^{t}v_{n}^{t}\,dx/t$ is the standard scalar product in $L^{2}((0,t);\mathbb{R})$ rescaled by a factor $1/t$. More precisely, since the goal is to solve an approximation problem on $I_{t}$, in order to define integrability we will consider the Lebesgue measure $\mathcal{L}^{1}$ on $\mathbb{R}$ restricted to $I_{t}$ and we will define $\forall t>0$ the rescaled measure $\mathcal{L}^{1}_{t}$ such that $\mathcal{L}^{1}_{t}(A)=\mathcal{L}^{1}(A)/t$ for all $A\subset I_{t}$. [^14] Notice that $\mathcal{L}^{1}_{t}$ is a probability measure on $I_{t}$ since beside being a well defined measure we also have that $\mathcal{L}^{1}_{t}(I_{t})=1$. One we have this measure we can define the Hilbert space $L^{2}_{\mathcal{L}^{1}_{t}}(I_{t};\mathbb{R})$ which is exactly the space of square $\mathcal{L}^{1}_{t}$-integrable, real-valued functions. So it is natural to require that the method that we will develop works on functions $u\colon\overline{\mathbb{R}}_{+}\to\mathbb{R}$ such that for all $t>0$ $u|_{I_{t}}\in L^{2}_{\mathcal{L}^{1}_{t}}(I_{t};\mathbb{R})$. The approximation problem then can be stated as the problem of finding a solution to the following minimization problem,151515 This problem has always a unique solution since the subspace $V^{t}_{N}$ is finite dimensional and hence it is closed
@@
where $V^{t}_{N}\subset L^{2}_{\mathcal{L}^{1}_{t}}(I_{t};\mathbb{R})$ is a finite, $N$-dimensional subspace that we assume to be spanned by $N$ orthonormal basis functions $v^{t}_{0},\dots,v^{t}_{N-1}$; i.e., $V^{t}_{N}:=\mathop{\rm span}\nolimits\{v^{t}_{0},\dots,v^{t}_{N-1}\}$. Here othornormality as usual means that $(v^{t}_{i},v^{t}_{j})_{t}=\delta_{ij}$ for all $i,j=0,\dots,N-1$ where $\delta_{ij}$ is the usual Kronecker delta and $(\cdot,\cdot)_{t}$ is the standard scalar product in $L^{2}_{\mathcal{L}^{1}_{t}}(I_{t};\mathbb{R})$, that is $(f,g)_{t}:=\int_{I_{t}}fg\,d\mathcal{L}^{1}_{t}\equiv(\int_{I_{t}}fg\,dx)/t$ being $dx$ the usual notation for the Lebesgue measure. In general the solution to the problem in Eq. (36) (see [147]) is given by
@@
[…]
@@
where $\Phi$ is the transition function that depends on the precise recurrent architecture, $u_{t}$ is the input of the net at time $t$ and $c_{t-1}$ are the coefficients of the online approximation of the function $f_{t}=\pi(x_{t})$, where $\pi\colon\mathbb{R}^{d_{\text{s}}}\to\mathbb{R}$ is a projection of the state onto the real line (which is necessary, since these online approximation methods work on scalar functions). The leap from this hybrid model, where the state of the recurrence is enriched with an online approximation of the state itself, to Deep State-Space Models has been proposed in [40], which analyzes Linear State Space Layers (LSSLs) in comparison to other deep learning models that are used to process sequences, and in [25], which refines such models to address computational limitations of the former model. In Linear State Space Layers (Figure 13), the main idea is to use a linear continuous-time expression to model the update of the state itself,
@@
The continuous time model described by Eq. (41) is typically discretized in order to be numerically implemented. Different discretization techniques can be applied, the more direct being the explicit (or forward) Euler method,161616 Technically this is a mixed scheme since the input is computed at the step $t+1$, however we call it explicit since it is so with respect to the state variable.
@@
This is what is commonly referred to as the convolutional form of the linear state space model in Eq. (44) – see Figure 14-top. Now, going back to LSSL [40], the matrix $A$ is represented with a suitable matrix factorization, i.e., $A=P(D+T^{-1}Q)$, with $D$, $P$ and $Q$ diagonal matrices and $T$ tridiagonal. The HiPPO matrix $A(t)$ defined in Eq. (39) admits such factorization (see Appendix E.2 of [40]). In this way, matrix $A$ is guaranteed to be quasiseparable, a property that is presented as desirable both for handling long dependencies and for enabling efficient matrix-vector multiplication. As a common practice in deep learning, several LSSLs can be stacked together, each layer receiving as input the output of the previous one. This is possible since the input and output have the same dimensionality. The main problem with this architecture is the computational cost; indeed (see [25]), it requires $\mathcal{O}({d_{\text{s}}}^{2}L)$ operations and it scales as $\mathcal{O}({d_{\text{s}}}L)$ for what concerns memory in order to compute the input-output mapping in Eq. (44). In order to overcome this precise limitation the S4 model [25] has been introduced to condition the structure of the matrix $A$. There is a large set of works that were published in the last few years along this line of research, and that, starting from S4, we describe in the following. Refer to Figure 15 for an overview.
@@
where $\text{diag}(\lambda_{1},\dots,\lambda_{{d_{\text{s}}}})$ is a diagonal matrix, $\lambda_{i}\in\mathbb{C}$ for every $i$ and $P,Q\in\mathbb{C}^{{d_{\text{s}}}\times 1}$. In Eq. (46), the $\dagger$} operation denotes the conjugate transpose and the term $PQ^{\dagger}$ is usually referred to as low-rank correction. With this particular choice for $A$, the computation of the recursion in Eq. (41) has complexity $\tilde{\mathcal{O}}({d_{\text{s}}}+L)$. For discretizing the dynamics in Eq. (41) and Eq. (42), the bilinear transform with discretization step $\tau$ is applied, leading to the already introduced,
@@
From SISO to MIMO: S5.   The Simplified Structured State Space Sequence Model (S5) [42] is the first Deep SSM to be parameterized leveraging the concept of multiple input multiple output (MIMO) systems, for simplifying the architectural components and enhancing computations. This means that the full input vector $u\in\mathbb{R}^{d_{\text{in}}}$ is fed into a single bigger MIMO system of Eq. (48), instead of $d_{\text{in}}$} SISO scalar smaller subsystems, by stacking the matrices $A^{\tau},\,B^{\tau},\,C$, used in S4 and S4D. S5 inherits the S4D parametrization of the matrix $A$ (i.e., it is a diagonal matrix), while it can be discretized applying both bilinear (as done in S4, see Eq. (47)) and exact (as done in S4D, see Eq. (50)) discretizations. The MIMO structure and the diagonal parameterization allows for parallel computation of the individual output components via a parallel scan algorithm (see Appendix H of [42]). The parallel scan algorithm (see Figure 14-bottom) offers a way to parallelize the computations of a sequence of elements of a semigroup $(S,\bullet)$ generated by a recurrence relation of the the form $s_{i+1}=s_{i}\bullet c_{i}$, where $(c_{i})_{i=1}^{L}$ is a given sequence of elements of SS.171717 We recall that a semigroup consist of a set $S$ together with an associative operation $\bullet$. This approach can be directly applied to the computation of a linear recurrence of the form in Eq. (44) with the following choices,
@@
Figure 15: Timeline showing the chronological development of deep state-space models—Overview of the organization of Section 4.
SSMs in Language Modeling: H3.   The Hungry Hungry Hippo (H3) [151] model is a novel approach to leverage SSMs for language modeling, aiming to address the limitations of previous SSMs in tasks like Associative Recall and Induction Heads (see Section 7) compared to attention-based models. H3 draws inspiration from linear attention, which assumes a specific form for the similarity metric used in attention calculations. It stacks two discrete SSMs: one with a shift matrix (i.e., a local convolution) to remember past tokens and one with a diagonal matrix to retain state over the entire sequence. The key innovation lies in introducing gates (i.e., multiplicative interactions) between the outputs of these SSMs and projections of the input . Combined with the shift matrix, these gates enable H3 to compare tokens across the sequence. The shift SSM identifies specific events (like the presence of a key token), while the diagonal SSM stores and retrieves associated information (like the corresponding value token). H3 has a time complexity of $\mathcal{O}(L\log(L))$ for a sequence of length $L$, making it asymptotically more efficient than traditional attention, which has a complexity of $\mathcal{O}(L^{2})$.
@@
where $C^{\tau}_{t}=C_{t}$, $D^{\tau}_{t}=D_{t}$ and $W_{\tau}\in\mathbb{R}^{1\times{d_{\text{in}}}}$. The model is structured in a MIMO manner (as S5, LRU and S6) and the dynamic matrix $A$ is initialized with $\lambda_{j}=-j$, for every $j=1,\dots,{d_{\text{s}}}$, ensuring that the eigenvalues lie in the negative halfplane. Since $\tau_{t}$ is time-varying, the eigenvalues of $A^{\tau}_{t}$ have an initialization that is input-dependent. The time-varying representation presents computational challenges, despite being more expressive. The authors provide an efficient implementation of the time-varying dynamics in Eq. (48), presenting a variation of the parallel scan and exploiting it both at inference and training time. S6 introduces an innovative way of pre-processing the input, called Mamba, which relies on both linear and non-linear maps. The input enters the recurrence through a linear projection, followed by a causal convolution. Additionally, it passes through a linear projection followed by a SiLU nonlinearity before entering the gating function for post-processing. The gating function is inspired by previous models, i.e., H3 and GAU. An architecture close to Mamba is the Gated State Space (GSS) layer [133], again inspired by GAU. GSS resembles Mamba but includes additional projections. The key difference is that GSS’s projection reduces the model dimension to decrease the state size of the SSM, whereas Mamba expands the model dimension to increase the state size.
@@
for $i=0,\dots,{d_{\text{in}}}-1$, where $h^{j}_{t}$ are implicit long convolution filters learned by shallow feed-forward neural networks and $R\in\mathbb{R}^{{d_{\text{in}}}\times{d_{\text{in}}}}$ is an output projection that mixes channels across the sequence length. This approach decouples filter length from parameter cost, providing advantages over explicit parameterization. The number of recurrent steps determines the complexity of the operator , and they can be represented as a decomposition of data-controlled matrices. These matrices dynamically adapt based on the input data, similar to how attention mechanisms compute a weighted sum over input elements. Instead of explicitly computing the full data-controlled matrix, Hyena leverages fast convolution algorithms, particularly Fast Fourier Transform (FFT)-based convolutions, to achieve subquadratic time complexity. Moreover, unlike some models that restrict the receptive field, it allows for interactions between any elements in the sequence through its long convolutions, enabling it to capture long-range dependencies effectively.
In subsequent work [155], some improvements to further enhance the efficiency of Long Convolution Sequence Models (LCSMs), including Hyena, have been introduced : the LaughingHyena distilling. It focuses specifically on improving the inference stage of these models, particularly in auto-regressive generation tasks. LaughingHyena distillation aims to represent each convolutional filter of a pre-trained LCSM as an SSM, with the smallest state dimension such that it approximates the original filter without significant loss of accuracy. To achieve this goal, it utilizes a method called modal interpolation, which provides coefficients for the numerator and denominator of a rational function that minimizes the difference between the original filter and the approximating SSM transfer functions. Through these coefficients, it is possible to define the matrices $A$, $B$, and $C$ which characterize the SSM. This distillation procedure is then followed by two steps: ($i$) pre-filling and ($ii$) a recurrent update rule. Pre-filling involves computing the state to start generating new tokens when a length-$L$ prompt is fed to the model during auto-regressive generation, exploiting the denominator of the approximate transfer function. The recurrent update rule for the complex state is defined as follows
@@
Theoretical Foundations of SSMs.  Performances achieved by SSMs are remarkable, thus inspiring several research efforts to understand both their expressive capabilities and the connections to existing popular technologies (such as attention), with which they share many features but have been commonly developed in isolation. Orvieto et al. [158] theoretically show that combining MLPs with either real or complex linear diagonal recurrences (such as in S4, Mamba, etc.) enables highly precise approximation of regular causal sequence-to-sequence maps. The proof is based on the fact that the linear RNN provides a lossless encoding of the input sequence, and the MLP conducts nonlinear processing on this encoding. While real diagonal linear recurrences are sufficient for achieving universality, employing complex eigenvalues near the unit disk, a strategy that has shown empirical success in S4, significantly improves the ability of recurrent models to store information. Cirone et al. [159] leverage tools from Rough Path Theory, and provide theoretical grounding for the fact that, when random linear recurrences are enhanced with simple input-controlled transitions (selectivity mechanism), the hidden state is demonstrably a low-dimensional projection of a mathematical construct known as the signature of the input. This signature captures nonlinear interactions between tokens across different timescales. Other recent works focus on the connections and differences between SSMs and other sequence processing models [32, 160], as long as to their links with control theory [161]. Deep SSMs such as S4, Mamba and H3 demonstrate that stacking linear recurrences with carefully chosen spectra can match or exceed attention on long contexts while retaining $\mathcal{O}(L)$ time and memory. Key themes include (i) diagonal, structured or low-rank transition matrices; (ii) convolutional or parallel-scan implementations for hardware efficiency; and (iii) theoretical guarantees that a linear core plus shallow non-linearity is a universal causal operator. These results confirm that recurrence—when stabilised—remains a first-class tool for sequence modelling.
The SSMs literature reveals that careful structuring of linear recurrence can recover scalability. Building on that insight, the community has revisited traditional RNNs themselves, refining gating, constraining weight geometry, or casting the dynamics in continuous time, to push their performance further. The following Section 5 reviews these architectural enhancements to “plain” recurrent networks.